A book with a mass of 0.56 kg is resting on a shelf. One of the supports holding the shelf is loose and the shelf suddenly tilts to one side..  If the angle  is 25.0o (measured from the horizontal), how fast would the book accelerate)



1
Expert's answer
2020-12-25T15:21:29-0500

Suppose that there is no friction. So,


"mg\sin\alpha=ma\to a=g\sin\alpha=9.8\cdot\sin25°=4.1(m/s^2)" . Answer